sorry, i didnt realize that the entire question didnt post. solve quadratic equation by factoring and using the zero-product principle:
28x^2+20x=0

Answer by user_dude2008(1862)

28x^2+20x=0

4x(7x+5)=0

4x=0 or 7x+5=0

x=0 or x=-5/7